# Decoupling Understanding from Reasoning via Problem Space Mapping for Small-scale Model Reasoning

### v1

@@ Abstract
Despite recent advances in the reasoning capabilities of Large Language Models (LLMs), improving the reasoning ability of Small Language Models (SLMs, e.g., $\leq$ 1.5B) remains challenging. A key obstacle lies in the complexity and variability of natural language: essentially equivalent problems often appear in diverse surface forms, often obscured by redundant or distracting details. This imposes a dual burden on SLMs: they must first extract the core problem from complex linguistic input, and then perform reasoning based on that understanding. The resulting vast and noisy problem space hinders optimization, particularly for models with limited capacity. To address this, we propose a new framework that decouples understanding from reasoning by mapping natural language problems into a canonical problem space—a semantically simplified yet expressive domain. This enables SLMs to focus on reasoning over standardized inputs, free from linguistic variability. Within this framework, we introduce DURIT (Decoupled Understanding from Reasoning via Iterative Training), a three-step algorithm that iteratively: (1) mapping natural language problems via reinforcement learning, (2) aligns reasoning trajectories through self-distillation, and (3) trains reasoning policies in the problem space. The mapper and reasoner are co-trained in an alternating loop throughout this process. Experiments show that DURIT substantially improves SLMs’ performance on both in-domain and out-of-domain mathematical and logical reasoning tasks. Beyond improving reasoning capabilities, DURIT also improves the robustness of reasoning, validating decoupling understanding from reasoning as an effective strategy for strengthening SLMs.
@@ Introduction
In this paper, we propose a general framework that maps natural language problems into a more abstract, low-dimensional, and semantically canonical problem space, effectively reducing the complexity of the original problem space. Within this space, models can learn and reason more efficiently. We instantiate this framework with a novel three-step alternating training algorithm: (1) a problem-space mapper is trained using RL and implicit templates to map natural language problems into standardized, low-dimensional forms; (2) self-distillation transfers this mapping capability into a SLM; and (3) a reasoning model is trained via RL to operate within the problem space. The mapper and the SLM are optimized in an alternating fashion, enabling iterative improvement in reasoning ability. To validate the effectiveness of DURIT, we conduct comprehensive empirical studies using models from the LLaMA (Grattafiori et al. 2024) and Qwen (Yang et al. 2025a) families, with parameter sizes ranging from 0.5B to 1.5B. Even when trained solely on the GSM8K (Cobbe et al. 2021) mathematical reasoning dataset, DURIT achieves significant gains across a range of in-domain and out-of-domain datasets, including those focused on mathematical and logical reasoning, and demonstrates strong generalization capabilities. Distinct from traditional CoT-based approaches, DURIT introduces a new paradigm that improves reasoning by reducing the complexity of the problem space. Our main contributions are summarized as follows:
@@
Based on this formulation, we propose a unified framework (Figure 1) that leverages a dedicated problem space mapper to project natural language questions into a standardized representation. By clustering fundamentally similar problems, this mapping reduces the exploration space and improves both sample and exploration efficiency during SLM training. As the model advances within this space, its ability to solve more complex problems increases, gradually shifting the underlying problem distribution. To adapt, our framework adopts an iterative training paradigm that alternates between updating the problem space mapper and refining the reasoning model, enabling their co-evolution. Reducing the problem space dimensionality enhances exploration and speeds up convergence. To analyze this effect, we consider a simplified Upper Confidence Bound setting and show through the following theorem that the regret bound decreases with problem space dimensionality.
@@
Let $\mathcal{Q}$ be a finite set of natural language problems, viewed as distinct states $s$, and let $A$ denote the set of candidate responses. At each round $t\in\{1,\dots,T\}$, a SLM observes a problem $s_{t}\in\mathcal{Q}$, selects an action $a_{t}\in A$, and receives a reward $r(s_{t},a_{t})$. Suppose learning is performed via a state-wise Upper Confidence Bound (UCB) algorithm in a contextual bandit setting. Then, the total regret after $T$ rounds is bounded by
@@ Methods
To decouple understanding from reasoning, a problem space mapper $M$ is instantiated as an LLM that maps natural language questions into a standardized problem space. While explicit templates enforce standardization, they are labor-intensive and may impede comprehension by SLMs. To balance standardization and flexibility, an implicit template mechanism is proposed, using a codebook to softly guide the output style of $M$. The mapping aims to 1) improve SLMs’ understanding and 2) reduce the complexity of the problem space. To facilitate this, we cluster the training data based on fundamental question similarity using k-Nearest Neighbors (kNN) (Cover and Hart 1967) over representations $\mathbf{z}_{i}$ encoded from each question $Q_{i}$, its description, and answer via model $M$. As no ground-truth labels exist, we adopt GRPO (Shao et al. 2024) to optimize $M$ based on the average correctness $r_{\text{acc}}$ of frozen SLM’s responses to mapped problem $Q_{i}^{\prime}$. To prevent $M$ from solving the problem directly, we apply a cheating penalty $r_{\text{cheating}}$ if $Q_{i}^{\prime}$ includes solution-specific terms (e.g., keywords like “answer value”) not present in $Q_{i}$. The total reward is:
@@
However, RL alone cannot sufficiently enforce standardization. To simplify the problem space, implicit templates conditioned on cluster labels $t_{i}$ are introduced. Specifically, a codebook $C$ of $n$ implicit template tokens $\{T_{1},\ldots,T_{n}\}$ and corresponding query keys $\{k_{1},\ldots,k_{n}\}$ is constructed, with both $\{T_{i}\}$ and $\{k_{i}\}$ randomly initialized. During training, for each problem $Q_{i}$, the template token $T_{t_{i}}$ is selected and concatenated with the original input as $x_{i}=[Q_{i};T_{t_{i}}]$, guiding $M$ to produce the mapped question $Q^{\prime}_{i}$. To encourage alignment between $Q^{\prime}_{i}$ and its assigned template, we define a template similarity loss based on the InfoNCE (He et al. 2020) objective:
$\mathcal{L}_{\text{template-sim}}=-\log\frac{\exp\left(\frac{\langle\mathbf{z}_{i},\mathbf{T_{i}}\rangle}{\tau}\right)}{\sum_{j=1}^{n}\exp\left(\frac{\langle\mathbf{z}_{i},\mathbf{T}_{j}\rangle}{\tau}\right)},$ (3)
where $\mathbf{z}_{i}$ is the normalized representation of the mapped problem $Q^{\prime}_{i}$ and $\tau$ is a temperature hyperparameter. At inference, with $t_{i}$ unavailable, the best-matching implicit template is selected via cosine similarity between the input question embedding $\mathbf{q}_{i}$ (approximated by averaging the word embeddings) and learned template query keys. A key similarity loss is introduced to facilitate key learning:
@@
Gradients from $\mathbf{q}_{i}$ are detached to prevent interference with the training of $M$, and only the template keys are updated. The overall loss function jointly optimizes the mapping policy and template-based constraints:
@@
$\mathcal{L}_{i}=\frac{1}{l}\sum_{k=1}^{l}\left[(1-\lambda)(-\log p_{s}(x_{i}^{k}))+\lambda\,\mathrm{KL}\big{(}p_{t}(x_{i}^{k})\parallel p_{s}(x_{i}^{k})\big{)}\right],$ (7)
@@
After distilling the transformation capability into the SLM, the model is further trained via RL to explore and reason directly in the original problem space, leveraging its internalized understanding. Specifically, we fine-tune the SLM on the original training dataset $\mathcal{D}_{0}$ using the GRPO algorithm, with answer correctness serving as the reward signal. As the reasoning model improves, its ability to interpret and generalize evolves, potentially altering the optimal structure of the problem space. To accommodate this, the problem space mapper $M$ and the reasoning model $R$ are trained in an iterative loop, enabling continual refinement of the problem space and progressive enhancement of reasoning capabilities.
@@
We conduct all experiments by training models exclusively on the GSM8K (Cobbe et al. 2021) dataset. For evaluation, we consider both in-domain (IND) and out-of-domain (OOD) settings. Specifically, we use GSM8K-Platinum (Vendrow et al. 2025) for IND assessment, and for evaluating mathematical reasoning generalization under distribution shift, we adopt MAWPS (Koncel-Kedziorski et al. 2016), SVAMP (Patel, Bhattamishra, and Goyal 2021), MATH500 (Hendrycks et al. 2021), and GAOKAO (Zhang et al. 2024) datasets. Additionally, to assess broader reasoning capabilities, we evaluate models on logical reasoning with the LogiQA (Liu et al. 2020) dataset. This comprehensive evaluation setup enables us to systematically analyze the impact of DURIT on the overall reasoning ability of SLMs across diverse domains.
@@
We compare our approach against several representative baselines across four categories: (1) CoT Distillation: including the Std-CoT (Magister et al. 2023), and similar to the second stage of DURIT, we additionally introduce STaR (Zelikman et al. 2024), where for each question, we sample $N$ CoT responses using the SLM and filter out the correct ones based on the final answer for fine-tuning; (2) Prompt Optimization: PRewrite (Kong et al. 2024), which similarly employs RL to optimize prompts; (3) RL-Based Methods: GRPO (Shao et al. 2024); and (4) Knowledge Distillation: we adopt Vanilla-KD (Muralidharan et al. 2024), which requires online inference of the teacher LM to obtain the token-level output distributions. In our setup, the mapper model serves as the teacher. Consistent with prior work (Sheng, Li, and Zeng 2025), we adopt answer accuracy as the primary evaluation metric. Further baseline details are available in the Appendix B.3B.
@@
To evaluate the generalization capability of DURIT, we test different base models, including recent strong instruction-following and reasoning-oriented models such as Qwen2.5-0.5B-Instruct (Yang et al. 2025a) and Llama3.2-1B-Instruct (Grattafiori et al. 2024). For the mapper model, we use Qwen2.5-3B-Instruct for the Qwen family, and Llama3.2-3B-Instruct for the Llama family to ensure architectural homogeneity. The codebook contains 32 implicit templates, with loss coefficients $\alpha_{1}=1\mathrm{e}{-3}$ for template query keys and $\alpha_{2}=1\mathrm{e}{-2}$ for tokens. Training is conducted in three steps: Step I runs for 1 epoch, Step II for 5 epochs, and Step III for 3 epochs. All experiments are carried out on 2 A100 GPUs with 40GB memory. For inference, we employ greedy decoding without vLLM (Kwon et al. 2023) acceleration. Additional implementation details can be found in Appendix B.3B. We also evaluate our approach on the Qwen2.5-1.5B-Instruct model (Yang et al. 2025a); detailed results and illustrative examples are provided in the Appendix E.1E and FF.
@@
Table 2: Comparison of different methods on Qwen2.5-0.5B-Instruct and Llama3.2-1B-Instruct. DURIT is trained with a single iteration. Orig: original test set; Symb: gsm-symbolic; $\Delta\%$: relative performance drop from Orig to Symb. Bold and underline indicate best and second-best results in each group.
@@
Figure 3: Comparative analysis of DURIT performance under different training data, mapper models, and prompt designs.
To evaluate the impact of different iterative training datasets on DURIT, we conducted a second iteration using GSM8K, MATH, and a filtered version of DeepScaleR (Luo et al. 2025b), following the first iteration based on GSM8K. As illustrated in Figure 3(a), the second iteration yields more pronounced improvements when the training data differs from that used in the first iteration. This finding highlights DURIT’s strength in decoupling understanding from reasoning, enabling it to leverage complementary data sources effectively. Moreover, training with more diverse datasets consistently improves overall performance and enhances the model’s reasoning capabilities. Detailed descriptions of the datasets are provided in the Appendix B.2B.
@@
To evaluate the impact of different models used as mappers on the performance of DURIT, we fix the reasoning SLM as Qwen2.5-0.5B-Instruct and perform one iteration of DURIT updates with various mapper models. Specifically, we experiment with Qwen2.5-3B/1.5B/0.5B-Instruct and Llama3.2-3B-Instruct, aiming to assess the effect of both model scale and model family on DURIT. Given the relatively weak instruction-following ability of Qwen2.5-0.5B-Instruct, we warm-start it with 200 mapper data generated by Qwen2.5-3B-Instruct to improve its initial alignment behavior. The results, presented in Figure 3(b), indicate that mappers from the same model family generally outperform those from other families, and that performance tends to improve slightly with increasing mapper model size. However, overall performance differences across mappers remain marginal, demonstrating that DURIT is robust to the choice of mapper. Even with a lightweight mapper like Qwen2.5-0.5B-Instruct, DURIT achieves strong performance without relying on external larger models.
@@
To assess the impact of mapper prompt design on the performance of DURIT, we experiment with three distinct prompt formulations (see Appendix B.3B), using Qwen2.5-3B-Instruct as the mapper and Qwen2.5-0.5B-Instruct as the reasoning SLM. Each configuration is trained for one iteration under identical conditions. As shown in Figure 3(c), performance varies slightly across prompts: more explicit and standardized prompts tend to yield better results. Nevertheless, all prompt variants lead to strong performance, indicating that DURIT is robust to variations in mapper prompt design.
@@
We conduct ablation studies using the Qwen2.5-0.5B-Instruct model to evaluate the contribution of each component in DURIT. For Step I, we assess the impact of the implicit template constraint by replacing mapper with PRewrite (w/o tem), while keeping the subsequent procedures in Step II and Step III of DURIT unchanged. For Step II, we examine the role of self-distillation by removing it and retaining only the SFT loss (w/o sd) and the necessity of SFT by removing sft loss(w/o sft). For Step III, we investigate the effect of removing GRPO training (w/o grpo). As shown in Table 3, ablating any single component leads to performance degradation. w/o tem disrupts the standardization of the problem space, resulting in less compact representations and lower exploration efficiency. w/o sd has minimal impact on in-domain performance but substantially impairs out-of-domain generalization, underscoring the role of self-distillation in reducing the comprehension burden and enhancing robustness. w/o sft may impose excessive disruption on the model’s inherent reasoning mechanisms and simultaneously expose it to biased or incorrect reasoning patterns in the mapped questions, potentially resulting in further performance degradation. Finally, w/o grpo consistently reduces accuracy, confirming the necessity of RL to strengthen reasoning after self-distillation.
@@
Table 3: Ablation study on Qwen2.5-0.5B-Instruct across six benchmark datasets. We report accuracy (%) on GSM8K (abbreviated as GSM), MAWPS (MW), SVAMP (SV), MATH500 (MT), GAOKAO (GK), and LogiQA (LQ). All variants are trained with a single iteration of DURIT.
@@
Table 4: Average 5-NN distance of Qwen2.5-0.5B-Instruct final hidden states under different input questions. Lower values indicate tighter local clustering.

### v2

@@ Abstract
Despite recent advances in the reasoning capabilities of Large Language Models (LLMs), improving the reasoning ability of Small Language Models (SLMs, e.g., up to 1.5B parameters) remains challenging. A key obstacle lies in the complexity and variability of natural language: essentially equivalent problems often appear in diverse surface forms, often obscured by redundant or distracting details. This imposes a dual burden on SLMs: they must first extract the core problem from complex linguistic input, and then perform reasoning based on that understanding. The resulting vast and noisy problem space hinders optimization, particularly for models with limited capacity. To address this, we propose a new framework that decouples understanding from reasoning by mapping natural language problems into a canonical problem space-a semantically simplified yet expressive domain. This enables SLMs to focus on reasoning over standardized inputs, free from linguistic variability. Within this framework, we introduce DURIT (Decoupled Understanding from Reasoning via Iterative Training), a three-step algorithm that iteratively: (1) mapping natural language problems via reinforcement learning, (2) aligns reasoning trajectories through self-distillation, and (3) trains reasoning policies in the problem space. The mapper and reasoner are co-trained in an alternating loop throughout this process. Experiments show that DURIT substantially improves SLMs’ performance on both in-domain and out-of-domain mathematical and logical reasoning tasks. Beyond improving reasoning capabilities, DURIT also improves the robustness of reasoning, validating decoupling understanding from reasoning as an effective strategy for strengthening SLMs.
Codehttps://github.com/monster476/DURIT
@@ Introduction
In this paper, we propose a general framework that maps natural language problems into a more abstract, low-dimensional, and semantically canonical space, effectively reducing the complexity of the original problem space. Within this space, models can learn and reason more efficiently. We instantiate this framework with a novel three-step alternating training algorithm: (1) a problem-space mapper is trained using RL and implicit templates to map natural language problems into standardized, low-dimensional forms; (2) self-distillation transfers this mapping capability into a SLM; and (3) a reasoning model is trained via RL to operate within the problem space. The mapper and the SLM are optimized in an alternating fashion, enabling iterative improvement in reasoning ability. To validate the effectiveness of DURIT, we conduct comprehensive empirical studies using models from the LLaMA (Grattafiori et al. 2024) and Qwen (Yang et al. 2025a) families, with parameter sizes ranging from 0.5B to 1.5B. Even when trained solely on the GSM8K (Cobbe et al. 2021) dataset, DURIT achieves significant gains across a range of in-domain and out-of-domain datasets, including those focused on mathematical and logical reasoning, and demonstrates strong generalization capabilities. Unlike traditional CoT-based approaches, DURIT introduces a paradigm that enhances reasoning by compressing the problem space. Our main contributions are summarized as follows:
@@
Based on this formulation, we propose a unified framework (Figure 1) that leverages a dedicated problem space mapper to project natural language questions into a standardized representation. By clustering fundamentally similar problems, this mapping reduces the exploration space and improves both sample and exploration efficiency during SLM training. As the model advances within this space, its ability to solve more complex problems increases, gradually shifting the underlying problem distribution. To adapt, our framework adopts an iterative training paradigm that alternates between updating the problem space mapper and refining the reasoning model, enabling their co-evolution. Reducing the problem space dimensionality enhances exploration and speeds up convergence. Follow (Cui et al. 2025), we model the problem using a bandit setting and apply a simplified Upper Confidence Bound, showing that the regret bound decreases with problem space dimensionality through the following theorem.
@@
Let $\mathcal{Q}$ be a finite set of natural language problems, viewed as distinct states $s$, and let $A$ denote the set of candidate responses. At each round $t\in\{1,\dots,T\}$, a SLM observes a problem $s_{t}\in\mathcal{Q}$, selects an action $a_{t}\in A$, and receives a reward $r(s_{t},a_{t})$. Suppose learning is performed via a state-wise Upper Confidence Bound (UCB) algorithm in a contextual bandit setting. Then, in the state-independent worst case, the total regret after $T$ rounds is bounded by
@@ Methods
To decouple understanding from reasoning, a problem space mapper $M$ is instantiated as an LLM that maps natural language questions into a standardized problem space. While explicit templates enforce standardization, they are labor-intensive and may impede comprehension by SLMs. To balance standardization and flexibility, an implicit template mechanism is proposed, using a codebook to softly guide the output style of $M$. The mapping aims to 1) improve SLMs’ understanding and 2) reduce the complexity of the problem space. To facilitate this, we cluster the training data based on fundamental question similarity using k-Nearest Neighbors (kNN) over representations $\mathbf{z}_{i}$ encoded from each question $Q_{i}$, its description, and answer via model $M$. As no ground-truth labels exist, we adopt GRPO (Shao et al. 2024) to optimize $M$ based on the average correctness $r_{\text{acc}}$ of frozen SLM’s responses to mapped problem $Q_{i}^{\prime}$. To prevent $M$ from solving the problem directly, we apply a cheating penalty $r_{\text{cheating}}$ if $Q_{i}^{\prime}$ includes solution-specific terms (e.g., keywords like “answer value”) not present in $Q_{i}$. The total reward is:
@@
However, RL alone cannot sufficiently enforce standardization. To simplify the problem space, implicit templates conditioned on cluster labels $t_{i}$ are introduced. Specifically, a codebook $C$ of $n$ implicit template tokens $\{T_{1},\ldots,T_{n}\}$ and corresponding query keys $\{k_{1},\ldots,k_{n}\}$ is constructed, with both $\{T_{i}\}$ and $\{k_{i}\}$ randomly initialized parameters and optimized by loss. During training, for each problem $Q_{i}$, the template token $T_{t_{i}}$ is selected and concatenated with the original input as $x_{i}=[Q_{i};T_{t_{i}}]$, guiding $M$ to produce the mapped question $Q^{\prime}_{i}$. To encourage alignment between $Q^{\prime}_{i}$ and its assigned template, we define a template similarity loss based on the InfoNCE (He et al. 2020) objective:
$\mathcal{L}_{\text{template-sim}}=-\log\frac{\exp\left(\frac{\langle\mathbf{z}_{i},\mathbf{T_{t_{i}}}\rangle}{\tau}\right)}{\sum_{j=1}^{n}\exp\left(\frac{\langle\mathbf{z}_{i},\mathbf{T}_{j}\rangle}{\tau}\right)},$ (3)
where $\langle\cdot,\cdot\rangle$ denotes the inner product, $\mathbf{z}_{i}$ is the normalized representation of the mapped problem $Q^{\prime}_{i}$ and $\tau$ is a temperature hyperparameter. At inference, with $t_{i}$ unavailable, the best-matching implicit template is selected via cosine similarity between the input question embedding $\mathbf{q}_{i}$ (both $\mathbf{z}_{i}$ and $\mathbf{q}_{i}$ are approximated by averaging the word embeddings) and learned template query keys. A key similarity loss is introduced to facilitate key learning:
@@
Gradients from $\mathbf{k}_{i}$ are detached to prevent interference with the training of $M$, and only the template keys are updated. The overall loss function jointly optimizes the mapping policy and template-based constraints:
@@
$\begin{split}\mathcal{L}_{i}&=\frac{1}{l}\sum_{k=1}^{l}\Big[(1-\lambda)\big(-\log p_{s}(x_{i}^{k})\big)\\ &\qquad\qquad+\ \lambda\,\mathrm{KL}\!\big(p_{t}(x_{i}^{k})\parallel p_{s}(x_{i}^{k})\big)\Big].\end{split}$ (7)
@@
After distilling the transformation capability into the SLM, the model is further trained via RL to explore and reason directly in the original problem space, leveraging its internalized understanding. Specifically, we fine-tune the SLM on the original training dataset $\mathcal{D}_{0}$ using the GRPO algorithm, with answer correctness serving as the reward signal. As the reasoning model improves, its ability to interpret and generalize evolves, potentially altering the optimal structure of the problem space. To accommodate this, the problem space mapper $M$ and the reasoning model $R$ are trained iteratively, enabling continual refinement of the problem space and progressive enhancement of reasoning capabilities.
@@
All experiments train models solely on GSM8K (Cobbe et al. 2021). Evaluation considers both in-domain (IND) and out-of-domain (OOD) settings: GSM8K-Platinum (Vendrow et al. 2025) for IND, and MAWPS (Koncel-Kedziorski et al. 2016), SVAMP (Patel, Bhattamishra, and Goyal 2021), MATH500 (Hendrycks et al. 2021), and GAOKAO (Zhang et al. 2024) for OOD mathematical reasoning. Broader reasoning is evaluated on LogiQA (Liu et al. 2020). This setup enables systematic analysis of DURIT’s impact on SLM reasoning across diverse domains.
@@
We compare DURIT against representative baselines in four categories: (1) CoT Distillation: including Std-CoT (Magister et al. 2023) and STaR (Zelikman et al. 2024), where $N$ CoT responses per question are sampled and correct ones are filtered for fine-tuning; (2) Prompt Optimization: PRewrite (Kong et al. 2024) using RL to optimize prompts; (3) RL-Based Methods: GRPO (Shao et al. 2024); and (4) Knowledge Distillation: Vanilla-KD (Muralidharan et al. 2024), which requires online teacher LM inference. In our setup, the mapper model serves as the teacher. Following prior work (Sheng, Li, and Zeng 2025), answer accuracy is the primary metric. Further baseline details are available in the Appendix B.3B.
@@
To evaluate the generalization capability of DURIT, we test different base models, including recent strong instruction-following and reasoning-oriented models such as Qwen2.5-0.5B-Instruct (Yang et al. 2025a) and Llama3.2-1B-Instruct (Grattafiori et al. 2024). For the mapper model, we use Qwen2.5-3B-Instruct for the Qwen family, and Llama3.2-3B-Instruct for the Llama family to ensure architectural homogeneity. The codebook contains 32 implicit templates, with loss coefficients $\alpha_{1}=1\mathrm{e}{-3}$ and $\alpha_{2}=1\mathrm{e}{-2}$. Training is conducted in three steps: Step I runs for 1 epoch, Step II for 5 epochs, and Step III for 3 epochs. All experiments are carried out on 2 A100 GPUs with 40GB memory. For inference, we employ greedy decoding without vLLM (Kwon et al. 2023) acceleration. Additional implementation details, as well as more experimental results, parameter analyses, and training time comparisons, can be found in Appendix E E.
@@
Table 2: Comparison of methods on Qwen2.5-0.5B-Instruct and Llama3.2-1B-Instruct. DURIT is trained with a single iteration. Orig: original test set; Symb: gsm-symbolic; $\Delta\%$: relative drop from Orig to Symb. Bold and underline indicate best and second-best results.
@@
Figure 3: DURIT performance across six benchmarks with varying training data, mapper models, and prompts: GSM (GSM8K), MW (MAWPS), SV (SVAMP), MT (MATH500), GK (GAOKAO), LQ (LogiQA), and average (Avg).
To evaluate the impact of iterative training datasets on DURIT, we conducted a second iteration using GSM8K, MATH, and a filtered DeepScaleR (Luo et al. 2025b), following the first iteration on GSM8K. As Figure 3(a) shows, improvements are more pronounced when the second-iteration data differs from the first. This demonstrates DURIT’s ability to decouple understanding from reasoning, effectively leveraging complementary data. Additionally, training with more diverse datasets consistently enhances overall performance and reasoning capabilities. Dataset details are provided in the Appendix B.2B.
@@
To evaluate the impact of different mappers on DURIT, we fix the reasoning SLM as Qwen2.5-0.5B-Instruct and perform one iteration of DURIT updates with various mappers (Qwen2.5-3B/1.5B/0.5B-Instruct and Llama3.2-3B-Instruct) to assess both model scale and family effects. Given the relatively weak instruction-following of Qwen2.5-0.5B-Instruct, we warm-start it with 200 mapper data from Qwen2.5-3B-Instruct to improve initial alignment. Results (Figure 3(b)) show that mappers within the same family generally outperform others, and performance slightly improves with larger model size. Overall differences are marginal, demonstrating DURIT’s robustness to mapper choice: even with a lightweight mapper like Qwen2.5-0.5B-Instruct, strong performance is achieved without relying on external larger models.
@@
To assess the impact of mapper prompt design on DURIT, we test three prompt formulations (see Appendix B.3B) using Qwen2.5-3B-Instruct as the mapper and Qwen2.5-0.5B-Instruct as the reasoning SLM, training each configuration for one iteration. As Figure 3(c) shows, performance varies slightly: more explicit, standardized prompts generally perform better. All variants achieve strong performance, indicating DURIT’s robustness to mapper prompt variations.
@@
We conduct ablation studies using the Qwen2.5-0.5B-Instruct model to evaluate the contribution of each component in DURIT. For Step I, we assess the impact of removing the implicit template constraint component (w/o tem), while keeping the subsequent procedures in Step II and Step III of DURIT unchanged. For Step II, we examine the role of self-distillation by removing it and retaining only the SFT loss (w/o sd) and the necessity of SFT by removing sft loss(w/o sft). For Step III, we investigate the effect of removing GRPO training (w/o grpo). As shown in Table 3, ablating any single component leads to performance degradation. w/o tem disrupts the standardization of the problem space, resulting in less compact representations and lower exploration efficiency. w/o sd has minimal impact on in-domain performance but substantially impairs out-of-domain generalization, underscoring the role of self-distillation in reducing the comprehension burden and enhancing robustness. w/o sft may impose excessive disruption on the model’s inherent reasoning mechanisms and simultaneously expose it to biased or incorrect reasoning patterns in the mapped questions, potentially resulting in further performance degradation. Finally, w/o grpo consistently reduces accuracy, confirming the necessity of RL to strengthen reasoning after self-distillation.
@@
Table 3: Ablation study of Qwen2.5-0.5B-Instruct on six benchmarks with a single DURIT iteration.
@@
Table 4: Average 5-NN distance of Qwen2.5-0.5B-Instruct final hidden states across different input questions. Lower values indicate tighter local clusters.